OpenStudy (anonymous):

Solve this system of linear equations. Separate the x- and y-values with a comma. 12x=54-6y -17x=-62-6y

OpenStudy (anonymous):

subtract the second equation from the first to get 12x-(-17x)=54-6y-(-62-6y) 29x=116 x=4 put this back to one of the above equations (I do the first) 12*4=54-6y 6y=6 y=1